# Evaluating Named Entity Recognition: Comparative Analysis of Mono- and Multilingual Transformer Models on Brazilian Corporate Earnings Call Transcriptions

### v1

@@ 5.2 Data annotation
To fine-tune the BERT models, it was necessary to assign an integer to each label. Hence, we compiled a list of the 23 labels in alphabetical order and created tags following the BIO scheme. For example, tags such as B-BALANCO_PATRIMONIAL and I-BALANCO_PATRIMONIAL were generated for the label BALANCO_PATRIMONIAL. Sequential numbers were then assigned to each tag, where 0 corresponds to “O”, 1 corresponds to B-BALANCO_PATRIMONIAL, 2 corresponds to I-BALANCO_PATRIMONIAL, and so forth. This list was saved in a JSON file and utilized in both the fine-tuning and evaluation steps.
6 Representation of the NER task as a text generation task
@@
The T5 model transforms text-based language problems into a text-to-text format Raffel2020 . Consequently, the token classification problem must be represented in a text-to-text format. In token classification, each token in a sentence is classified individually. For instance, a BERT model fine-tuned for NER processes a tokenized sentence and assigns a class to each token.
@@
Various approaches to using T5 for NER are found in the literature, all representing the token classification task as a text generation task. Carmo et al. Carmo2020 input sentences into their T5-based model (PTT5) preceded by the instruction “Recognize Entities” and the model outputs the same sentence, but annotated with entity labels enclosed in brackets immediately following each entity. Wang et al. Wang2022instructionner provide a T5 model with sentences preceded by an instruction and expect a list of sentences following the template “entity is a/an type”.
@@
In our approach, similar to Carmo et al. Carmo2020 , we input sentences into the model and expect the same sentence with entities and their labels. Unlike Carmo et al. and Wang et al. Wang2022instructionner , we do not use a preceding instruction because we fine-tuned our models for a single task. Wang et al. argue that their instruction helps the model identify labels and avoid omissions or additions. However, our testing revealed that this instruction led to longer inputs, increasing computational resource consumption during fine-tuning. Comparison of Precision, Recall, and macro F1-scores with and without the instruction favored omission of the instruction, resulting in inputs with fewer tokens, lower resource consumption, and higher metric values.
@@
For example, given $S$ as “John lives in New York” the model outputs $S^{\prime}$ as “[John—PERSON] lives in [New_York—LOCATION]”. Hence, rather than classifying individual tokens or subwords, the model generates sentences with annotated entities.
@@ 7.1 BERT-based models fine-tuning
After reviewing Table 4, it becomes apparent that the number of tokens remains consistent across models, as this calculation is model-independent. Particularly, mBERT produces a higher number of subwords compared to BERTimbau, with a ratio of 1.33 for mBERT and 1.19 for BERTimbau. This ratio plays a crucial role in determining model hyperparameters, as it helps estimate the number of tokens in the model input. Additionally, it aids in decision-making processes regarding potential truncation of inputs or padding methods.
@@
Upon analyzing the sentence lengths in terms of the number of subwords, we noted that the longest sentence in mBERT comprised 379 subwords, while in BERTimbau, the maximum was 324 subwords. Based on this observation, we concluded that neither of the BERT models would require truncation of any sentences. Consequently, we proceeded with the fine-tuning of both BERTimbau and mBERT, utilizing batch sizes of 16, 2 epochs, and a learning rate of $5\cdot 10^{-5}$ (the default value of Hugging Face Transformers TrainingArguments [^10] https://huggingface.co/docs/transformers/main_classes/trainer) for each model. These hyperparameter values align with those outlined by Devlin et al. Devlin2019 and are deemed suitable for the available infrastructure, which includes an NVIDIA T4 GPU with 15 GB of memory.
@@
Model GPU Mem Time Precision Recall F1
@@ 7.2 T5-based models fine-tuning
For T5-based models, each occurrence in every dataset underwent processing, involving the identification and annotation of each entity in the target sentence following our approach (cf. Section 6), and the tokenization of both sentences (input) and target (label). Ultimately, similar to BERT models (cf. Section 7.1), we obtained lists of “input_ids” and “labels” and saved the processed datasets in the Hugging Face DatasetDict structure for use in fine-tuning.
@@
Analyzing Table 6, we observed that the vocabulary size of mT5 is nearly 7.8 times larger than that of PTT5. Upon calculating the intersection between the vocabularies, we found that there are 14,459 subwords in common, which is nearly the same number as the intersection between BERTimbau and mBERT. Further analysis of the ratios revealed an increase from input to target sentences for both models due to tokens added by the annotation. Additionally, we confirmed that the maximum number of subwords in input sentences was 324, while in target sentences, it reached 509 subwords for PTT5. For mT5, the corresponding maximum numbers were 480 and 511 for input and target sentences, respectively. Consequently, we set the hyperparameters of the models to 512 to avoid truncation and conserve computational resources.
@@
Model Vocab Input Target
@@
Model GPU Mem Time Precision Recall F1
@@ 8 Evaluation
The fine-tuned models were employed for inference on the Test dataset using batches of size 16. The computational resources utilized during the inference step are presented in Fig. 3. In this figure, the models, along with the GPUs they were fine-tuned on, are arranged based on the memory consumed in GB (x-axis) and the time consumed in minutes (y-axis).
In Fig. 3, a significant disparity is evident between BERT and T5 models concerning the total time for inference. While BERT models required less than 5 minutes and nearly 5 GB of memory to classify the subwords of 5,794 sentences from the Test dataset, T5 models required over 15 minutes to generate the annotated sentences. Comparing memory usage, it’s notable that mT5 consumed nearly the same amount as BERTimbau and mBERT, whereas PTT5 consumed more than 13 GB.
Figure 3: Computational resources consumed during the inference step
Observing the BERT models, in Fig. 3, we can note that BERTimbau and mBERT models exhibited slight differences in memory and time consumed. While BERTimbau consumed 4.5 GB of memory and 2 minutes to classify the subwords, mBERT consumed 5.6 GB and 2.5 minutes. It is noteworthy that PTT5 required more computational resources. The model consumed 13.2 GB of memory and took 27 minutes to run on a T4 GPU, while it consumed 23.5 GB of memory and 20 minutes on an A100 GPU. In contrast, mT5 consumed fewer resources, utilizing 4.3 GB and 4.7 GB of memory and running for 24 and 18 minutes on T4 and A100 GPUs, respectively. The reason why PTT5 consumed nearly 10 GB more memory on the A100 GPU was not evident, and further investigation is required.
Fig. 4 depicts the Precision, Recall, and macro F1-score (F1) values in percentage (%), plotted using a logarithmic scale. Upon examining the performance of the models, we noted consistent values for these metrics across different GPUs for both PTT5 and mT5. Therefore, we only plotted the data for the T4 GPU for PTT5 and mT5.
@@
Analyzing Fig. 4, it is evident that BERTimbau achieved the highest Precision (99.41%) and macro F1-score (98.99%) among all models. Ranking the models by descending F1-score, we have BERTimbau (98.99%), mBERT (98.86%), mT5 (98.85%), and PTT5 (98.52%). Although mBERT’s F1-score is slightly lower than that of BERTimbau, it demonstrates a better balance between Precision and Recall compared to the other models. It is worth noting that the F1-scores of mBERT and mT5 are nearly identical, with values of 98.86% and 98.85%, respectively.
Comparing mono- and multilingual models, BERTimbau outperformed mBERT in terms of F1-score and computational resources required for fine-tuning and inference. Conversely, in the case of T5 models, mT5 exhibited superior performance compared to PTT5 when considering performance metrics. This difference may be attributed to the larger size of mT5, which has 80 million more parameters than PTT5. However, further investigation is needed to explore the specific characteristics of each architecture and pre-training process. Furthermore, the higher number of subwords of multilingual models seems to have no considerable impact on their performance when compared to monolingual models.
Upon examining the measurements per class of all models (A), we observed consistently high values overall. Generally, BERT-based models achieved an F1-score higher than 0.98, while T5-based models scored higher than 0.97. However, we noted that BERTimbau and mBERT obtained the lowest F1-scores in the class RESULTADO, with 0.8667 and 0.8462, respectively. Analyzing the results for PTT5 and mT5, we also observed a low value of 0.9565 for both models in the class RESULTADO. Additionally, PTT5 and mT5 achieved their lowest F1-scores (0.9504 and 0.9144) in the class YEAR.
@@
The nested entities and the number of entities are likely the cause of the low-performance scores for all models in the class RESULTADO. Nested entities are a particular challenge and were echoed by Shah et al. Shah2023 , who noted similar difficulties encountered by their model when handling nested entities.
After computing performance metrics, we proceeded to calculate metrics related to model errors, adapted from MUC-5 (cf. Section 4.4). For this task, we manually verified each classification by comparing the label assigned during the annotation step with the label assigned by the model. If the model was correct, we marked it with a “Yes”; otherwise, we used a “No”.
The results of this evaluation are presented in Table 8. In this table, we included: the metric abbreviation; an indication of whether the model assigned the correct label; and the corresponding values for each model. For example, BERTimbau had 10 classifications marked as Incorrect (INC), but upon manual verification, we found that the model was correct in 3 classifications, indicating that the annotations in the dataset were incorrect.
Table 8: Values of adapted MUC-5 metrics
@@
COR 30060 33114 16378 16445
MIS No 44 44 98 53
Yes 1 0 5 3
INC No 7 14 8 7
Yes 3 5 7 0
SPU No 7 25 4 92
Yes 97 95 108 63
Error per response fill 0.1919 0.2493 0.6623 0.9122
Undergeneration 0.1461 0.1326 0.5941 0.3211
Overgeneration 0.0232 0.0752 0.0242 0.5540
Substitution 0.0233 0.0423 0.0488 0.0425
Analyzing the error metrics in Table 8, we observe that BERTimbau exhibits lower values for key metrics. Specifically, in terms of “Error per response fill”, BERTimbau demonstrates correctness in more situations compared to the other models. Additionally, it excels in “Overgeneration”, generating proportionally fewer incorrect spurious classifications, and in “Substitution”, making fewer changes from correct to incorrect classifications. On the other hand, mBERT performs better in “Undergeneration”, missing fewer classifications than the other models.
We also observe in Table 8 that PTT5 generated more correct additional classifications (108) than the other models, followed by BERTimbau (97). According to MUC-5, the term “Spurious” is interpreted as an error because the model classified more tokens than those annotated in the dataset. However, it is important to note that we anticipated that the models, after fine-tuning, would be capable of identifying and classifying entities not annotated in the original dataset, resulting in additional correct predictions. Therefore, from our perspective, it is a positive scenario when a model generates correct additional classifications.
While manually analyzing the classifications generated by PTT5 and mT5, we observed instances of text repetition and alterations in the generated text. For example, out of the 92 spurious classifications by mT5, 76 were repetitions, indicating that the model generated 16 distinct “real” spurious classifications. We further categorized the incorrectly generated text as either “critical” or “non-critical”. An example of a critical error is the alteration of a numerical representation of money: from the original target [R\$_824,00|MONEY] to the generated [R\$_2824,00|MONEY]. In this case, the original entity was “824,00”, and the model erroneously generated “2824,00”, representing a difference of two thousand, corresponding to a number more than three times bigger. Non-critical errors are related to minor changes in the text that do not alter the sentence’s meaning.
Observing Figure 5, it is evident that PTT5 generated 460 sentences with errors — 294 with critical errors and 166 with non-critical errors. This accounts for nearly 8% of the Test dataset. In contrast, mT5 generated 201 sentences with errors, representing almost 3.5% of the dataset — 71 with critical errors and 130 with non-critical errors.
@@
Using the similarity ratios, we examined the frequency distribution of sentences categorized as critical or non-critical, along with their respective similarity ratios. For PTT5, among the 294 sentences with critical errors, an overwhelming majority (98%) exhibited a similarity ratio falling within the range of 0.96 to 1.0. In the case of mT5, a significant portion of both the 71 sentences with critical errors (94.4%) and the 130 sentences with non-critical errors (96.2%) fell within the similarity ratio range of 0.89 to 1.0.
@@
9 Final Considerations
In this study, we developed a novel dataset tailored for Named Entity Recognition (NER) tasks within the financial domain, specifically in Portuguese. Comprising 23 labels encompassing both generic and domain-specific categories, this dataset is the first of its kind. Extracted from 384 earnings call transcriptions of ten prominent Brazilian banks, it provides a comprehensive resource for training and evaluating NER models. Furthermore, we conducted performance evaluations on four Transformer-based models: BERTimbau and PTT5, which were pre-trained exclusively in Portuguese, alongside multilingual models mBERT and mT5. This comprehensive assessment sheds light on the efficacy of these models in capturing financial entities and demonstrates their applicability in real-world scenarios.
In the comparison between BERT models, BERTimbau outperformed mBERT in both the fine-tuning and evaluation phases. Despite achieving a slightly higher macro F1-score in the evaluation phase, BERTimbau demonstrated lower memory and time consumption compared to mBERT. While these differences may seem minor, they can become more significant when processing larger datasets. Therefore, in our context, choosing BERTimbau as the preferred model is feasible.
Within the T5 models, when working within constraints on computational resources for fine-tuning, PTT5 is a suitable choice due to its lower memory requirements for fine-tuning than mT5. However, given the difference in F1-scores between PTT5 and mT5, mT5 could be preferred as it requires less memory and processing time during evaluation. If additional computational resources are available for a short period, fine-tuning mT5 and subsequently using it in a more restricted infrastructure could be a viable option.
In evaluating error metrics adapted from MUC-5, it becomes evident that BERTimbau consistently outperforms mBERT across various metrics, except for “Undergeneration”, where mBERT exhibits a lower rate of missed classifications compared to BERTimbau. Similarly, PTT5 demonstrates superior performance over mT5 in metrics such as “Error per response fill” and “Overgeneration”, indicating fewer incorrect subword classifications and a reduced number of spurious outputs. However, mT5 exhibits a lower rate of missed classifications and demonstrates a slight advantage in maintaining label consistency between correct and incorrect classifications. Despite the models’ F1-score, when comparing MUC-5 metrics, the monolingual models consistently achieve better error metric values.

### v2

@@ 4.1.2 Data annotation
4.2 Models fine-tuning
We searched on the Hugging Face platform for models provided by research groups or companies, specifically focusing on those pre-trained either exclusively in Portuguese or on multilingual datasets. We identified BERTimbau and PTT5 as models pre-trained in Portuguese, and mBERT and mT5 as multilingual models. While reviewing the available versions on Hugging Face, we opted for versions recommended by the original authors and with a comparable number of parameters to facilitate result comparison. Additionally, we included the smallest versions due to infrastructure constraints, despite the expectation of better results with larger models [17, 18].
Table 3: Selected models
Architecture Model Language Version #Parameters
BERT BERTimbau Portuguese Base 110M
BERT mBERT Multilingual Base 110M
T5 PTT5 Portuguese Base 220M
T5 mT5 Multilingual Small 300M
We studied the models’ tokenizers and their inputs and outputs. After that, we developed scripts using Google Colaboratory (Colab) platform, Python v3.10, and Hugging Face libraries such as Transformers v4.35. We prepared the data developing different scripts for each architecture due to its characteristics. For example, while BERT-based models expect a sentence and a sequence of labels as inputs, T5-based models expect a sentence and an expected text (text that would be generated). For both architectures, we used the Beginning, Inside, and Outside (BIO) tagging scheme [33].
@@ 4.2 Models fine-tuning
To fine-tune the BERT models, it was necessary to assign an integer to each label. Hence, we sorted the 23 labels in alphabetical order and created tags following the BIO scheme. For example, tags such as B-BALANCO_PATRIMONIAL and I-BALANCO_PATRIMONIAL were generated for the label BALANCO_PATRIMONIAL. Sequential numbers were then assigned to each tag, where 0 corresponds to “O”, 1 corresponds to B-BALANCO_PATRIMONIAL, 2 corresponds to I-BALANCO_PATRIMONIAL, and so forth. This list was saved in a JSON file and utilized in both the fine-tuning and evaluation steps.
@@
The T5 model transforms text-based language problems into a text-to-text format [18]. Consequently, the token classification problem must be represented in a text-to-text format. In token classification, each token in a sentence is classified individually. For instance, a BERT model fine-tuned for NER processes a tokenized sentence and assigns a class to each token.
@@
Various approaches to using T5 for NER are found in the literature, all representing the token classification task as a text generation task. Carmo et al. [23] input sentences into their T5-based model (PTT5) preceded by the instruction “Recognize Entities” and the model outputs the same sentence, but annotated with entity labels enclosed in brackets immediately following each entity. Wang et al. [20] provide a T5 model with sentences preceded by an instruction and expect a list of sentences following the template “entity is a/an type”.
@@
In our approach, similar to Carmo et al. [23], we input sentences into the model and expect the same sentence with entities and their labels. Unlike Carmo et al. [23] and Wang et al. [20], we do not use a preceding instruction because we fine-tuned our models for a single task. Wang et al. [20] argue that their instruction helps the model identify labels and avoid omissions or additions. However, our testing revealed that this instruction led to longer inputs, increasing computational resource consumption during fine-tuning. Comparison of Precision, Recall, and macro F1-scores with and without the instruction favored omission of the instruction, resulting in inputs with fewer tokens, lower resource consumption, and higher metric values.
@@
For example, given $S$ as “John lives in New York”, the model outputs $S^{\prime}$ as “[John—PERSON] lives in [New_York—LOCATION]”. Hence, rather than classifying individual tokens or subwords, the model generates sentences with annotated entities.
4.3 Evaluation of the models’ results
We evaluated the models considering various aspects, including computational resource requirements for both fine-tuning and inference, performance metrics such as Precision, Recall, and macro F1-score, and error analysis based on the metrics from the Fifth Message Understanding Conference (MUC-5) [34].
To calculate MUC-5 metrics, it’s imperative to compare the labels assigned by the model with those assigned during the annotation step. The outcome is then classified into one of the following categories [34]:
  • 1.
  • 2.
  • 3.
  • 4.
We calculated Primary and Secondary metrics based on the categories, where lower values indicate better performance. The Primary metric “Error per response fill”, adapted from MUC-5 [34], is expressed by Eq. 2.
$Error\ per\ response\ fill=\frac{(INC+MIS+SPU)}{(COR+INC+MIS+SPU)}$ (2)
$Undergeneration=\frac{MIS}{COR+INC+MIS}$ (3)
$Overgeneration=\frac{SPU}{COR+INC+SPU}$ (4)
$Substitution=\frac{INC}{COR+INC}$ (5)
We also verified the statistical differences among the models using the Wilcoxon test and Friedman’s test, followed by the Nemenyi test [35, 36]. For Friedman’s test, since we have a small number of subsets and classifiers, we calculated both ${\chi}^{2}$ (Q) and F statistics [35, 36]. Our null hypothesis ($H_{0}$) for the tests was that there is no statistical difference among the models at a significance level of 5% ($\alpha=0.05$), while our alternative hypothesis ($H_{1}$) was that differences do exist among the models. To perform the tests, we split the test dataset into five subsets, each of which was evaluated with every model. Based on the collected values, we calculated the statistical tests using the packages Pingouin v0.5.4 [37] and SciPy v1.11.4 [38].
@@ 5.1 BERT-based models fine-tuning
After reviewing Table 4, it becomes apparent that the number of tokens remains consistent across models, as this calculation is model-independent. Particularly, mBERT produces a higher number of subwords compared to BERTimbau, with a ratio of 1.33 for mBERT and 1.19 for BERTimbau. This ratio plays a crucial role in determining model hyperparameters, as it helps estimate the number of tokens in the model’s input. Additionally, it aids in decision-making processes regarding potential truncation of inputs or padding methods.
@@
Upon analyzing the sentence lengths in terms of the number of subwords, we noted that the longest sentence in mBERT comprised 379 subwords, while in BERTimbau, the maximum was 324 subwords. Based on this observation, we concluded that neither of the BERT models would require truncation of any sentences. Consequently, we proceeded with the fine-tuning of both BERTimbau and mBERT, utilizing batch sizes of 16, 2 epochs, and a learning rate of $5\cdot 10^{-5}$ (the default value of Hugging Face Transformers TrainingArguments [^11] https://huggingface.co/docs/transformers/main_classes/trainer) for each model. These hyperparameter values align with those outlined by Devlin et al. [17] and are deemed suitable for the available infrastructure, which includes an NVIDIA T4 GPU with 15 GB of memory.
@@
Model GPU Memory Time Precision Recall F1
@@ 5.2 T5-based models fine-tuning
For T5-based models, each occurrence in every dataset underwent processing, involving the identification and annotation of each entity in the target sentence following our approach (cf. Section 4.2), and the tokenization of both sentences (input) and target (label). Ultimately, similar to BERT models (cf. Section 5.1), we obtained lists of “input_ids” and “labels” and saved the processed datasets in the Hugging Face DatasetDict structure for use in fine-tuning.
@@
Analyzing Table 6, we observed that the vocabulary size of mT5 is nearly 7.8 times larger than that of PTT5. Upon calculating the intersection between the vocabularies, we found that there are 14,459 subwords in common, which is nearly the same number as the intersection between BERTimbau and mBERT. Further analysis of the ratios revealed an increase from input to target sentences for both models due to tokens added by the annotation. Additionally, we confirmed that the maximum number of subwords in input sentences was 324, while in target sentences, it reached 509 subwords for PTT5. For mT5, the corresponding maximum numbers were 480 and 511 for input and target sentences, respectively. Consequently, we set the hyperparameters of the models to 512 to avoid truncation and save computational resources.
@@
Model Vocab. Input Target
@@
Model GPU Memory Time Precision Recall F1
@@ 6 Evaluation
The fine-tuned models were employed for inference on each subset of the Test dataset using batches of size 16. The mean and standard deviation (std) of the time utilized during the inference step are presented in Fig. 4. We can observe, in Fig. 4, a significant disparity between BERT and T5 models concerning the mean time for inference. While BERTimbau required a mean of 16.04 seconds ($\pm$ 3.58 s) to classify the subwords of the five subsets of the Test dataset, the PTT5 model required 28 minutes (an average of 338.14 $\pm$ 43.57 seconds) to generate the annotated sentences. Regarding memory usage, we observed that BERTimbau consumed 3.7 GB of memory, while mBERT and mT5 consumed 4.4 GB each. In contrast, PTT5 consumed 12.4 GB of memory, which is nearly three times the amount consumed by the other models.
Figure 4: Average $\pm$ standard deviation of inference time per model (in seconds)
Fig. 5 depicts the mean and standard deviation of Precision, Recall, and macro F1-score (F1) values in percentage (%). Upon examining the performance of the models, we noted consistent values for these metrics across different GPUs for both PTT5 and mT5. Therefore, we only plotted the data for the T4 GPU for PTT5 and mT5.
@@
Analyzing Fig. 5, it is evident that BERTimbau achieved the highest average Precision (99.44%) and macro F1-score (98.99%) among all models, with the lowest standard deviations (0.64% and 0.48%, respectively), indicating more stable performance across evaluations. When comparing mono- and multilingual models, BERTimbau outperformed mBERT in terms of F1-score and the computational resources required for fine-tuning and inference. Conversely, among the T5 models, mT5 exhibited superior performance compared to PTT5 in terms of performance metrics. This difference may be attributed to the larger size of mT5, which has 80 million more parameters than PTT5. However, further investigation is needed to explore the specific characteristics of each architecture and pre-training process. Furthermore, the higher number of subwords in multilingual models seems to have no considerable impact on their performance when compared to monolingual models.
Despite the slight differences among the models, Friedman’s test indicated a statistically significant difference (p-value $<$ 0.05) only for Precision (Q = 8.28, p-value = 0.04; F = 4.92, p-value = 0.02). Performing the Nemenyi test for Precision, we observed a statistically significant difference only between BERTimbau and PTT5 (p-value = 0.03). Analyzing Wilcoxon results, we observed that there is no statistically significant difference at the 5% level between each pair of models for each metric.
Upon examining the measurements per class of all models (A), we observed consistently high values overall. However, we noted that BERTimbau and mBERT obtained the lowest F1-scores in the class RESULTADO, with 0.8878 and 0.8667, respectively. Analyzing the results for PTT5 and mT5, we also observed a low value of 0.9333 for both models in the class RESULTADO. Additionally, PTT5 and mT5 achieved their lowest F1-scores (0.9505 and 0.9353) in the class YEAR.
@@
The nested entities and the number of entities are likely the cause of the low-performance scores for all models in the class RESULTADO. Nested entities are a particular challenge and were echoed by Shah et al. [15], who noted similar difficulties encountered by their model when handling nested entities.
After computing the performance metrics, we compared the labels assigned during the annotation step with those assigned by the model, following MUC-5 standards [34]. Additionally, we manually verified each classification, marking them with “Yes” when the model was correct and with “No” when the model was incorrect. The results of this evaluation are presented in Table 8. This table includes the class abbreviation, an indication of whether the model assigned the correct label, and the corresponding values for each model. For example, BERTimbau had 48 classifications marked as Missing (MIS), but upon manual verification, we found that the model was correct in 2 classifications, indicating that the annotations in the dataset were incorrect.
Table 8: Results of the manual verification
@@
COR 30058 33114 16362 16418
MIS No 46 44 108 61
Yes 2 0 8 7
INC No 9 14 12 8
Yes 0 0 0 0
SPU No 12 25 9 97
Yes 83 95 115 70
We observe in Table 8 that PTT5 generated more correct additional classifications (115) than the other models, followed by mBERT (95). According to MUC-5, the term “Spurious” is interpreted as an error because the model classified more tokens than those annotated in the dataset. However, it is important to note that we anticipated that the models, after fine-tuning, would be capable of identifying and classifying entities not annotated in the original dataset, resulting in additional correct predictions. Therefore, from our perspective, it is a positive scenario when a model generates correct additional classifications.
While manually analyzing the classifications generated by PTT5 and mT5, we observed instances of text repetition and alterations in the generated text. We further categorized the incorrectly generated text as either “critical” or “non-critical”. Non-critical errors are related to minor changes in the text that do not alter the sentence’s meaning, while critical errors involve changes in numerical values, words, and text repetition. An example of a critical error is the alteration of a numerical representation of money: from the original target [R\$_824,00|MONEY] to the generated [R\$_8924,00|MONEY]. In this case, the original entity was “824,00”, and the model erroneously generated “8924,00”, representing a nearly tenfold increase. Other examples include:
  • 1.
    In the entity “Receita Federal” (the Brazilian Federal Revenue Service) the model changed “Federal” to “fixa” (fixed) generating “receita fixa” (fixed revenue);
  • 2.
    In a sentence ending with the segment “[R\$_10,8_bilhões|MONEY] para acho que está em torno de [R\$_9,8|MONEY]” ([R\$_10.8_billion|MONEY] to I think it’s around [R\$9.8|MONEY]), the model altered “10,8” to “8,8” and repeated the entire segment eight times; and
  • 3.
    In the segment “prejuízo foram [0,08%|PERCENTUAL]” (losses were
    [0.08%|PERCENTAGE]), “0.08%” was changed to “0.88%”, an increase of eleven times.
Observing Figure 6, it is evident that PTT5 generated 345 sentences with errors — 299 with critical errors and 46 with non-critical errors. This accounts for nearly 6% of the Test dataset. In contrast, mT5 generated 130 sentences with errors, representing almost 2.3% of the dataset — 77 with critical errors and 53 with non-critical errors.
@@
Using the similarity ratios, we examined the frequency distribution of sentences categorized as critical or non-critical, along with their respective similarity ratios. For PTT5, among the 299 sentences with critical errors, an overwhelming majority (98%) exhibited a similarity ratio in the range of 0.93 to 1.0, while 96% of the sentences with non-critical errors had a similarity ratio between 0.97 and 1.0. In the case of mT5, out of the 77 sentences with critical errors, 70 (91%) had a similarity ratio from 0.91 to 1.0, while among the 53 sentences with non-critical errors, 43 (87%) had a similarity ratio ranging from 0.99 to 1.0.
@@
Based on the data presented in Table 8, we calculated MUC-5 metrics for the five subsets of the Test dataset. Table 9 presents the mean values of the metrics with their standard deviation. Analyzing these metrics, we observe that BERTimbau exhibits lower values for key metrics. Specifically, in terms of “Error per response fill”, BERTimbau demonstrates correctness in more situations compared to the other models. Additionally, it excels in “Overgeneration”, generating proportionally fewer incorrect spurious classifications, and in “Substitution”, making fewer changes from correct to incorrect classifications. On the other hand, mBERT performs better in “Undergeneration”, missing fewer classifications than the other models.
Table 9: Mean $\pm$ standard deviation values of adapted MUC-5 metrics
Metric BERTimbau mBERT PTT5 mT5
Error per response fill 0.2212 ($\pm$0.08) 0.2503 ($\pm$0.10) 0.7775 ($\pm$0.14) 0.9893 ($\pm$0.88)
Overgeneration 0.0397 ($\pm$0.02) 0.0754 ($\pm$0.03) 0.0544 ($\pm$0.05) 0.5790 ($\pm$0.79)
Substitution 0.0299 ($\pm$0.02) 0.0422 ($\pm$0.04) 0.0730 ($\pm$0.05) 0.0484 ($\pm$0.05)
Undergeneration 0.1522 ($\pm$0.07) 0.1334 ($\pm$0.07) 0.6563 ($\pm$0.14) 0.3696 ($\pm$0.16)
Regarding the Friedman’s test for the metrics in Table 9, we found a statistically significant difference at the 5% level for “Error per response fill” (Q = 12.12, p-value = 0.007; F = 16.83, p-value = 0.000337), “Undergeneration” (Q = 12.84, p-value = 0.005; F = 23.78, p-value = 0.000077), and “Substitution” (Q = 7.70, p-value = 0.0527; F = 4.21, p-value = 0.038). We calculated the Nemenyi test for these metrics and observed a statistically significant difference in “Error per response fill” between BERTimbau and PTT5 (p-value = 0.036). For “Undergeneration”, the statistically significant difference appears between mBERT and PTT5 (p-value = 0.003). Despite Friedman’s test indicating a statistically significant difference in F-statistic for “Substitution”, the Nemenyi test only showed a p-value of 0.068 between BERTimbau and PTT5.
After calculating the Wilcoxon test for each metric, we observed p-values of 0.062 or higher. Therefore, we accept the null hypothesis: there is no statistically significant difference between the pairs of models.
Based on the statistical tests, we can conclude that BERTimbau exhibited superior performance among the models. We also observed that there are no statistically significant differences between mono and multi-lingual models (BERTimbau versus mBERT and PTT5 versus mT5).
7 Conclusions
In this study, we evaluated four Transformer-based models: BERTimbau and PTT5, which were exclusively pre-trained in Portuguese, along with the multilingual models mBERT and mT5, for the financial Named Entity Recognition (NER) task. We introduced BraFiNER, a novel dataset specifically designed for NER tasks in the Portuguese financial domain. This dataset includes 23 labels covering both generic and domain-specific categories, making it the first of its kind.
Based on our evaluation, BERTimbau emerges as the preferred model due to its superior performance in both metrics and computational efficiency during inference. Statistical tests comparing mono- and multi-lingual models revealed no significant differences between them.